Octane contains only C and H. Combustion of 30g of octane produced 92.63 grams of CO2 and 42.63 grams H2O. The molar mass of octane is 114g. Calculate the empirical formula and molecular formula of octane.
1
Expert's answer
2020-12-25T05:11:32-0500

Any alkane on combustion will lead to CO2 and H2O formation.


No. of moles of CO2 formed = 92.63/44 = 2.10 moles

No. of moles of H2O formed = 42.63/18 = 2.36 moles

Moles of octane taken = 30/114 = 0.263 moles


so on taking ration of moles we get for 1 mole of octane we get 8 moles of carbon dioxide and

9 moles of water.


so total number of carbon in octane is 8 and hydrogen is 18.

Molecular formula is C8H18 and empirical formula is C4H9